the molar enthalpy of combustion of ethanol (c2H5OH) is -1366.8kj/mol what is the enthalpy change when 10.00 grams of ethanol is combusted


1
Expert's answer
2021-02-26T05:59:41-0500

n(C2H5OH)=m/M=10/46 = 0.2174 mol

H= (-1366.8)*0.2174 = -297.14 kJ

Answer: -297.14 kJ